A manchine with an efficiency 80% is used to raise a body of mass 75kg through a vertical height of 3m in 30s calculate the power input taking g=10ms-2
1
Expert's answer
2020-01-13T11:02:43-0500
"\rm The\: efficiency=\frac{Output\:power}{Input\:power}""\eta=\frac{mgh/t}{P}"

So, the input power


"P=\frac{mgh}{\eta t}=\frac{75\times 10\times 3}{0.8\times 30}=93.75\:\rm W"

Answer: "93.75\:\rm W"